OpenStudy (anonymous):

Which shows the values in increasing order? fifty-five percent, 5 for every 8, four-fifths, zero point seven zero point seven, four-fifths, 5 for every 8, fifty-five percent fifty-five percent, 5 for every 8, zero point seven, four-fifths zero point seven, four-fifths, fifty-five percent, 5 for every 8

OpenStudy (anonymous):

C because .55, .625, .7, .8

OpenStudy (anonymous):

55 % = .55 5 four 8 = 5/8 =.625 an last .7 four fifth = 4/5 = 0.8 so .55 , .625 , 0.7 , 0.8 hence C